For all parts of this question, assume the following: The CAPM holds. The riskless rate of return is 5%. The market portfolioplease show all work
Answer #1

The CAPM model shows that the points (return and stdv) which are below the capital market line are in infeasible reason. This means no investor, be it risk-taking or risk-neutral, won't invest in such portfolios.

return capital market line 15% risk free rates% not a good investment 4% 20% 30% standard dev

If a risk free asset is giving a return of 5%, then no one would go for an asset with 30% stdv (risky asset) to get 4% return. Hence, Linda is right.

2. Out of 35000 of available funds, 25000 (71.43%) are invested in Walmart and 28.57% are invested in tesla.

Expected return = W1*R1 +W2*R2 where W1 and W2 are the weights and R1 and R2 are the expected returns from each stocks.

hence, the expected return of the portfolio = 0.7143*5% + 0.2857*20%= 9.2858%

portfolio variance = (W1S1)^2 + (W2S2)^2 + 2*W1W2S1S2Cor, where S1 and S2 are stdv of portfolio and Cor is the correlation between these stocks

stdv of portfolio =( (0.7128*0.12)^2 + (0.2857*0.35)^2 + 2*0.7128*0.2857*0.12*0.35*0.2)^0.5 = 14.4%

Return Capital Market Line • Belchers current portfolio 407 14.4% 20% Stev

We can see from the figure that Bob's investment is not optimal as it lies below the market line.

If he wants to retain the same stdv, we need to find corresponding expected return on Capital market line, which is 12% return.

12% >= W1'*5% + W2'*20%

W1'= 1- W2'

12% = 5% - 5%*W2' +W2'*20%

W2 = 0.466 = 16333

Hence, he should invest 16333 in Tesla and remaining in Walmart